Compute the formula- weight concentration of NH3 in a solution in which it is known to be 3.0 percent ionized.



1
Expert's answer
2021-03-24T01:05:41-0400

Molar Mass of NH3 = 17.031g/ mol

formula- weight concentration = Mass of the solute/ volume of the solution

= Ionized NH3 > NH3 + 3% H

= 17.031 + 3% of 1.00784 = 17.031 + 0.0302352

= 17.0612352 g

1 g = 1 ml = 17.0612352 ml

= 17.031/17.0612351

= 0.9982g/ml